OpenStudy (anonymous):

Find the slope of line passing through the points (4,8) and (6, -2).

OpenStudy (anonymous):

do you know the formula for finding the slope?

OpenStudy (anonymous):

\[\frac{-2-8}{6-4}=\frac{-10}{2}=-5\]